find the equation of the parabola with its focus at (-4,7) and directrix y=1
Answer by lwsshak3(11628)

find the equation of the parabola with its focus at (-4,7) and directrix y=1
This is a parabola that opens upwards.
Its standard form of equation: (x-h)^2=4p(y-k), (h,k)=(x,y) coordinates of the vertex
For given parabola:
axis of symmetry: x=-4
x-coordinate of vertex=-4
y-coordinate of vertex=4 (half way between directrix and focus on axis of symmetry)=(7+1)/2=4
vertex: (-4,4)
p=3 (distance from vertex to focus or directrix on axis of symmetry
4p=12
Equation:
(x+4)^2=12(y-4)